COUPLING OF THE C-H STRETCH TO LARGE-AMPLITUDE TORSION AND INVERSION MOTIONS: COMPARISON OF CH_{3}CH_{2}^{.}, CH_{3}OH_{2}^{+} AND CH_{3}NH_{2}

\underline{RAM S. BHATTA}, Department of Polymer Science and Department of Chemistry, The University of Akron; DAVID S. PERRY, Department of Chemistry, The University of Akron, OH 44325.

In each of the title molecules, torsional and inversion tunneling occurs between six equivalent minima. Coupling of these degrees of freedom to the CH stretch occurs via variation of the C-H stretching force constants as a function of the torsional (α) and inversion (τ) angles. Maps of the couplings have been computed at the MP2/6-311++G(3df,2p) level. Both the single bond CH stretch force constants and the bilinear couplings between CH bonds are presented as a function of α and τ. Although the torsional barriers differ by more than a factor of 20, the torsion-inversion-vibration coupling patterns are very similar for CH_{3}NH_{2} and CH_{3}CH_{2}^{.}. On the other hand, the torsion-inversion-vibration coupling in the charged species CH_{3}OH_{2}^{+} is much weaker.